OpenStudy (anonymous):

Solve x – 3y = -17 using the elimination method. -x + 8y = 52

OpenStudy (rsadhvika):

you have the opposite coeffecients for x already

OpenStudy (rsadhvika):

simply add the equations vertically

OpenStudy (anonymous):

by adding both the equations x will be cancelled and you'l get y=7

OpenStudy (rsadhvika):

x – 3y = -17 -x + 8y = 52 ----------------

OpenStudy (rsadhvika):

add them and see wat u get

OpenStudy (anonymous):

now substituting y=1 we get x=-14

OpenStudy (anonymous):

sorry y= 7

OpenStudy (anonymous):

and x=4